A certain rain cloud at an altitude of 2.07 km contains 3.20 ✕ 107 kg of water vapor. How long would it take for a 2.55-kW pump to raise the amount of water from Earth's surface to the cloud's position?
Answer #1

power = work/time
And,
work = force x distance = m g x h = 3.2x10^7kg x 9.8m/s/s x 2070 = 64915.2 * 10^7 J
Now,
2.55kW means 2550J/s are generated by this pump; therefore, it will take a time of :
t = 64915.2x10^7J/2550J/s = 25.457 x 10^7 sec.